# Beyond A Single AI Cluster: A Survey of Decentralized LLM Training

### v2

@@ Abstract
The emergence of large language models (LLMs) has revolutionized AI development, yet the resource demands beyond a single cluster or even datacenter, limiting accessibility to well-resourced organizations. Decentralized training has emerged as a promising paradigm to leverage dispersed resources across clusters, datacenters and regions, offering the potential to democratize LLM development for broader communities. As the first comprehensive exploration of this emerging field, we present decentralized LLM training as a resource-driven paradigm and categorize existing efforts into community-driven and organizational approaches. We further clarify this through: (1) a comparison with related paradigms, (2) a characterization of decentralized resources, and (3) a taxonomy of recent advancements. We also provide up-to-date case studies and outline future directions to advance research in decentralized LLM training.
@@
Haotian Dong1thanks:   Equal Contribution., Jingyan Jiang2111More details about parallel strategies of distributed training are presented in Appendix A, Rongwei Lu1, Jiajun Luo1, Jiajun Song1, Bowen Li3, Ying Shen4, Zhi Wang1thanks:   Corresponding Author. 1Tsinghua University, 2Shenzhen Technology University, 3Beihang University, 4China Central Depository & Clearing Co., Ltd. donght24@mails.tsinghua.edu.cn, jiangjingyan@sztu.edu.cn wangzhi@sz.tsinghua.edu.cn
@@ 1 Introduction
The rapid advancement of LLMs has shown remarkable progress across various domains DeepSeek-AI et al. (2025b). With model scales expanding from GPT-3’s Brown et al. (2020) 175 billion to DeepSeek-R1’s DeepSeek-AI et al. (2025a) 660 billion parameters, the computing resource demands of training LLMs have increased dramatically Jiang et al. (2024).
However, this exponential growth in computational requirements poses significant challenges. For individual researchers and small laboratories with limited resources, the demands are particularly prohibitive. Even for well-resourced organizations, confining LLM training to a single AI cluster faces challenges like: geographically distributed services required for latency optimization McMahan et al. (2017), inherent hardware bottlenecks limiting single-cluster scalability Athlur et al. (2022); Grattafiori et al. (2024), economic patterns requiring placement adaptation Liu et al. (2023), etc.
@@
Duan et al. (2024) Efficient LLM Training, Centralized Infrastructures
Khan et al. (2023) Decentralized Machine Learning, Geo-Distributed Machine Learning
Woisetschläger et al. (2024) Federated Learning, Efficient Foundation Model Training
Ours Decentralized LLM training, Decentralized Infrastructures
@@
This survey systematically investigates challenges and solutions for decentralized LLM training. Compared to surveys of other distributed paradigms, our paper centers around LLM training with decentralized resources, as shown in Table 1. Our target is to characterize the utilization of decentralized resources and analyze optimization methods in decentralized LLM training, thereby exploring potential research opportunities. Figure 1 illustrates the utilization paradigms of resources in decentralized LLM training and optimization objectives we categorize in this paper. To the best of our knowledge, our survey is the first to study recent advances in decentralized LLM training. Our work contains three distinct contributions:
@@
  • Up-to-date case studies and prospective future research directions. We compare two LLMs trained under different decentralized paradigms: one leveraging organizational resources and the other utilizing scattered, fragmented resources. We also outline future directions on decentralized resources scaling law and governance for further development.
Figure 1: Decentralized LLM training paradigms of both community-driven and organizational decentralization. Community-driven decentralization utilizes pooling resources from communities with independent entities. These pooled resources can include local servers from researchers, instances from cloud services or volunteer communities, etc. Organizational decentralization involves consolidating resources from multiple clusters or even multiple data centers managed by well-resourced organizations (e.g., tech giants and governments).
@@ 2.1 Geo-ML and FL
Geo-ML and FL both tackle challenges from data decentralization, yet each approach has its distinct focus. Geo-ML mainly tackles service latency and regulatory needs by optimizing training processes across datacenters, leveraging hierarchical network topologies (e.g., high-bandwidth local area networks (LANs) within datacenters and limited-bandwidth WANs between datacenters) to utilize decentralized data and computational resources efficiently Hsieh et al. (2017). In contrast, FL focuses more on privacy protection McMahan et al. (2017), employing either centralized parameter-server or decentralized methods Lalitha et al. (2019); Xing et al. (2021), with resources from decentralized clusters or edge devices.
@@ 2.2 De-ML
De-ML has a dual meaning in terms of decentralization. From the distributed strategy perspective, it mainly addresses communication bottlenecks inherent in parameter server architectures through peer-to-peer networking Hegedűs et al. (2019); Warnat-Herresthal et al. (2021). From the resource utilization perspective, it offers a cost-effective and flexible strategy that maximizes the utilization of geographically dispersed computing resources Yuan et al. (2022). While both De-ML and Geo-ML involve training architectures, they differ fundamentally in their approach — the former adopts peer-to-peer topology where nodes communicate directly, while the latter relies on hierarchical centralized architecture both across and within datacenters.
@@ 2.4 Decentralized LLM Training
Decentralized LLM training leverage a broader range of resources compared to traditional distributed training, yet it still employ similar parallel strategies. Data parallelism (DP) and pipeline parallelism (PP) are the primary distributed strategies  [^1] More details about parallel strategies of distributed training are presented in Appendix A for training LLMs with decentralized resources Yuan et al. (2022); Ryabinin et al. (2023). Tensor parallelism (TP), due to intensive communication requirements, is more constrained within both LANs and WANs.
@@ 3.1 Communication Constraint
In distributed training, computational devices need to communicate frequently. But decentralized resources have more limited communication capabilities compared to centralized ones.
Community-driven decentralized resources, combining with local servers and cloud instances, communicate through LANs or even WANs during model training with bandwidths under 10 Gb/s Azurespeed (2024). In contrast, organization-driven decentralized training can reach over 500 GB/s of bandwidth within a server by NVLink, or over 200 Gb/s within a cluster using InfiniBand. However, when the scale of resources expands across multiple geographically distributed clusters or datacenters, the bandwidth can drop to only hundreds of Mb/s Xiang et al. (2022). To alleviate communication bottlenecks, strategies such as gradient compression Lu et al. (2024b) and delayed aggregation Zhu et al. (2021) are often employed during distributed training.
@@ 3.3 Cost Volatility
However, reducing costs with decentralized resources can be even more challenging than with centralized setups. In community-driven decentralization, fragmented resources such as preemptive cloud instances are relatively cheaper, but their prices are highly volatile Lee and Son (2017). Additionally, resource instability can lead to re-computation and inter-node waiting, prolonging training process and increasing overall costs. For organizational decentralization, accurately modeling energy consumption and optimizing resource scheduling become increasingly complex, owing to the demand for massive resources and the inherently cross-datacenter nature of this paradigm Faiz et al. (2024); Choudhury et al. (2024).
@@ 4 Community-Driven Decentralization
Training LLMs under community-driven paradigm often encounters challenges, including hardware heterogeneity, limited communication bandwidth, resources instability, and economic uncertainty. This chapter focuses on these challenges and provides an in-depth discussion of related research efforts. We summarize and compare selected studies in Table 2. A more comprehensive paper list is presented in Figure 2 of Appendix B.
@@
DeDloc Diskin et al. (2021) DP with extremely large batches BERT (170M) 91 unique devices (RTX 2060, K80, etc.)
@@ 4.1 Communication Optimization
In PP strategy, gradients are exchanged within a stage, while both gradients and activations are transmitted across stages. Varuna Athlur et al. (2022) leverages rule-based policy to adjust PP depth based on the available GPU count to better accommodate bandwidth constraints. To allocating tasklets requiring high communication to computing units with faster connections, DTFM Yuan et al. (2022) uses a two-level approach: a balanced graph partitioning problem for DP within each stage, and a joint graph matching and traveling salesman problem for the entire PP process. More dynamically, SWARM Ryabinin et al. (2023) optimizes the PP process by enabling real-time adjustments during each iteration. In this strategy, each PP stage uses multiple candidate devices. When a device outperforms others, it processes inputs from multiple slower predecessors and distributes outputs to multiple slow successors, maximizing bandwidth utilization.
@@ 4.2 Heterogeneity Awareness
Bandwidth constraints necessitate control of communication volume in decentralized LLM training, while network heterogeneity introduces additional challenges in communication scheduling.
To deal with network heterogeneity, BACombo Jiang et al. (2020) uses a bandwidth-aware worker selection strategy, enabling efficient gradient splitting and scheduling across multiple connections. DTFM Yuan et al. (2022) employs a communication matrix to model bandwidth differences, and it solves the hierarchical optimization problem to minimize communication cost. In contrast to static scheduling, DeDLOC Diskin et al. (2021) dynamically adapts its strategy by switching between All-Reduce, parameter servers, and decentralized SGD based on real-time network conditions. To further optimize communication paths, NETSTORM Li et al. (2024) employs a multi-root adaptive synchronization topology to dynamically allocate tasks based on bandwidth and balancing communication loads with auxiliary paths. Additionally, hardware characteristics can be integrated into the scheduling process. Holmes Yang et al. (2024) automatically selects heterogeneous network interface cards (NICs, including InfiniBand, RoCE, Ethernet) for diverse distributed strategies (DP, PP and TP) across heterogeneous clusters.
@@
11footnotetext: https://developer.nvidia.com/cuda-toolkit
@@ Discussion.
The effective utilization of decentralized resources is fundamentally constrained by the ability to manage system heterogeneity. Current frameworks rarely account for cross-architecture collaboration (e.g., NVIDIA GPUs with CUDA [^2] https://www.hiascend.com/software/cann and Huawei Ascend GPUs with CANN33footnotemark: 3), which limits their potential. Furthermore, existing heterogeneity-aware strategies mainly focus on protocol or topology optimization, overlooking critical semantic aspects of LLM training dynamics, such as layer sensitivity and gradient rank Refael et al. (2025). Developing semantics-guided, architecture-agnostic frameworks may enhance scalability and resource efficiency in decentralized LLM training.
@@ 4.4 Economics
Additionally, several studies propose analytical cost models to guide deployment decisions. Strati et al. [2024] develops a training cost estimator that highlights the benefits of intra-region communication for improving throughput and minimizing expenses. Phalak et al. [2024] extends this to a multi-cloud, multi-geography scenario, incorporating serverless compute and VM selection into a unified model for performance-cost optimization.
@@ 5 Organizational Decentralization
In organizational paradigms, well-resourced organizations often pursue optimization objectives that extend beyond individual systems to include datacenter-level enhancements, such as energy efficiency, system throughput within and across datacenters. Related papers are summarized in Figure 3 of Appendix B.
@@ 5.1 Energy Efficiency
Training LLMs within datacenters entails substantial energy consumption, making carbon efficiency a critical concern for large organizations. For a single job, predicting the carbon footprint of current training methods for large models can help optimize training strategies, thereby reducing carbon emissions Faiz et al. (2024). Resource selection across datacenters also plays a vital role, enabling a trade-off between model performance and environmental impact Bian et al. (2024) At the datacenter level, implementing job scheduling strategies that are aware of carbon emissions can further decrease the overall carbon footprint Park et al. (2024); Song et al. (2024). In addition, digital twin technologies show great promise in supporting real-time monitoring and optimization of datacenter operations for improved engery efficiency during LLM training Wang et al. (2023).
@@ 5.2 System Throughput
When training large models across datacenters, it’s crucial to design resource scheduling policies that account for the characteristics of training processes, such as 4D parallelism and synchronization requirements, as well as the infrastructure characteristics including low-bandwidth inter-datacenter communication Fan et al. (2024); Hu et al. (2024). Traditional cross-datacenter schedulers often fall short, as they struggle to adapt to the dynamic and resource-intensive nature of LLM training workloads Huang et al. (2019b). To address these challenges, MAST Choudhury et al. (2024), a global scheduling system, effectively orchestrates the training of LLaMA3 Grattafiori et al. (2024) across 16,000 H100 GPUs, achieving both load balancing and fault tolerance across datacenters.
@@ 6 Case Study
This section examines two representative models, Llama-3 and INTELLECT-1, which we considered as examples of organizational and community-driven decentralization, respectively. Comparative configurations are listed in Table 4 of Appendix D.
@@
Llama-3, an open-source foundation model series by Meta, scales up to 405 billion parameters Grattafiori et al. (2024). Its training processes 15 T multilingual tokens, requiring FLOPs. Computationally, Llama-3 utilizes 16,000 H100 GPUs with a global-scale scheduler Choudhury et al. (2024) and 4D parallel training, achieving 38-43% Model Fractional Utilization (MFU). Storage is managed by the Tectonic distributed file system Pan et al. (2021), offering 240 PB capacity and optimized throughput to reduce GPU idle time. Networking leverages RoCE and InfiniBand within a three-layer topology Lee et al. (2024); Gangidi et al. (2024), enhanced by load balancing and congestion control for efficient communication across 24,000 GPUs. As a representative open-source LLM, Llama-3’s training leverages interconnected clusters and a global-scale scheduler, offering valuable insights for organizational decentralized LLM training paradigms.
@@
INTELLECT-1 Jaghouar et al. (2024a) is claimed to be the first 10-billion-parameter LLM model trained with decentralized resources. It employs hierarchical parameter aggregation and int8 quantization to minimize bandwidth usage, while VPN technology enhances stability in low-bandwidth networks. For fault tolerance, INTELLECT-1 utilizes ElasticDeviceMesh for node management and phased checkpoint storage to minimize training interruptions, with peer-to-peer transfer enabling rapid checkpoint recovery. By integrating FSDP2 and CPU offloading, it optimizes device memory utilization, further enhancing the system’s efficiency and scalability. This community-driven decentralized paradigm democratizes AI model development, preventing monopolization and fostering open-source innovation with decentralized resources.
7 Summary and Future Directions
7.1 Summary
In this survey, we classify decentralized LLM training into two paradigms based on resource utilization: community-driven decentralization and organizational decentralization. By analyzing the characteristics of decentralized resources, reviewing relevant optimization methods from the literature, and examining two model cases, we provide a systematic overview of the current development landscape in decentralized LLM training.
7.2 Future Directions
Scaling Law of Decentralized LLM Training.
The scaling law for centralized LLM training primarily focus on computational power, data volume, and model size to optimize training strategies Hoffmann et al. (2022); Grattafiori et al. (2024). In decentralized paradigms, however, the interplay between computational and network resources becomes significantly more complex, necessitating their inclusion in the scaling laws. For a given topology of computational resources with certain network bandwidth, there may exist an upper bound to the optimal model size. Beyond this limit, increasing resources or data within single node or local group may not result better model performance. Exploring the scaling law for decentralized LLM training is essential for effectively coordinating global decentralized resources and enhancing their utilization efficiency.
Decentralized Resources Governance.
Current decentralized LLM training primarily focuses on the utilization and management of resources for individual model training processes. As the community expands, effective governance of decentralized resources will become crucial for sustaining the development of decentralized LLM training. Challenges like communication bottlenecks, resource heterogeneity, and instability may stem from the resources themselves or arise from inefficient coordination at the resource abstraction layer of decentralized systems. Future research could prioritize developing mechanisms to optimize pricing and scheduling of decentralized resources in multi-tenant environments, thereby facilitating the proliferation of open-source models powered by decentralized infrastructure.
8 Limitations
This survey focuses on decentralized LLM training from a resource-driven perspective, but several limitations should be noted. First, important issues such as data distribution and privacy protection are not covered, though they diverge from the scope of our survey. Second, given the rapid advancements in LLM development, some recent developments may have been inadvertently overlooked despite our efforts to include more relevant research. Additionally, this paper does not introduce specific techniques tailored to different types of LLMs, such as reasoning models or vision language models, while these models are also crucial for the utilization of decentralized resources and the reduction of AI monopolization.
Appendix A Parallel Strategies in LLM training
The training of LLMs demands substantial computational resources and sophisticated parallelization strategies to achieve efficiency. To address the challenges posed by the massive scale of LLMs, researchers have developed multi-dimensional parallelization approaches, including Data Parallelism (DP), Tensor Parallelism (TP), Pipeline Parallelism (PP), and Expert Parallelism (EP). Among these, TP and PP can be regarded as specialized forms of Model Parallelism (MP). When these techniques are strategically combined, they collectively enhance throughput, reduce memory footprint, and optimize resource utilization. Additionally, Context Parallelism (CP) NVIDIA (2023) has emerged as a complementary strategy, which operates at the token level by slicing input sequences across devices, enabling scalable and efficient training of long-context LLMs. The former four parallel strategies have been implemented within decentralized LLM training, especially DP and PP. A detailed comparison of these four parallelism strategies is presented in Table 3.
Aspect Data Parallelism Pipeline Parallelism Tensor Parallelism Expert Parallelism
Parallelism Granularity Input samples Model stages + Micro-batches Intra-layer tensors Experts (specialized sub-models)
Model copy per device Full model Partitioned by stages Partitioned by tensor slices Partitioned by expert assignments
Communication Payload Full gradients Inter-stage activations Intra-layer activations and gradients Expert-specific activations and gradients
Computation Concurrency High (across batches) Medium (scheduled pipeline) High (fine-grained operations) High (parallel expert execution)
Memory Usage Redundant model copy Stage-specific model partition Partial tensor per device Partial expert per device
Scalability Good for large batch sizes Good for inter-node communication Good for intra-node communication Good for inter-node communication
Representative Framework PyTorch DDP Li et al. (2020) GPipe Huang et al. (2019a) Megatron-LM Narayanan et al. (2021) Switch Transformer Fedus et al. (2022)
Table 3: Comparison of 4 primary parallel strategies in LLM training.
Appendix B Literature Summary
Studies associated with community-driven decentralization and organizational decentralization are presented in Figure 2 and Figure 3, respectively.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=black, rounded corners, align=left, text centered, minimum width=4em, edge+=black, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=8.2em,font=,, where level=2text width=12em,font=,, [ Organizational
Decentralization, fill=blue!5, font=[ Energy Efficiency §5.1, fill=yellow!5 [ TowardDC Wang et al. (2023),
CAFE Bian et al. (2024),
LLMCarbon Faiz et al. (2024),
CAFTM Park et al. (2024),
CloudSimPer Song et al. (2024). , leaf, text width= 11em, font=] ] [ System Throughput §5.2, fill=yellow!5 [ Yugong Huang et al. (2019b),
TFAS Fan et al. (2024),
MAST Choudhury et al. (2024),
ACME Hu et al. (2024).
, leaf, text width= 11em, font=] ] ]
Appendix C Distributed Hash Tables (DHTs)
Additionally, DHTs enable robust coordination of decentralized communication and update mechanisms, even under high node volatility. Moshpit Ryabinin et al. (2021) use DHTs to dynamically form groups for gradient averaging, ensuring that partial updates are aggregated reliably despite frequent node failures. These approaches collectively highlight that DHTs can enhance the fault tolerance during decentralized LLM training.
Appendix D Supplementary of Case Study
Table 4 shows the configurations of two representative LLMs trained with organizational and community-driven decentralized resources, respectively.
@@ Appendix D Supplementary of Case Study
Effective Training Time 90% 83%
Tokens Taken 15 TB 1 TB

### v3

@@ Abstract
The emergence of large language models (LLMs) has revolutionized AI development, yet their resource demands beyond a single cluster or even datacenter, limiting accessibility to well-resourced organizations. Decentralized training has emerged as a promising paradigm to leverage dispersed resources across clusters, datacenters and even regions, offering the potential to democratize LLM development for broader communities. As the first comprehensive exploration of this emerging field, we present decentralized LLM training as a resource-driven paradigm and categorize existing efforts into community-driven and organizational approaches. We further clarify this through: (1) a comparison with related paradigms, (2) characterization of decentralized resources, and (3) a taxonomy of recent advancements. We also provide up-to-date case studies and outline future directions to advance research in decentralized LLM training.
@@
Haotian Dong1thanks:   Equal Contribution., Jingyan Jiang2111More details about parallel strategies of distributed training are presented in Appendix A, Rongwei Lu1, Jiajun Luo1, Jiajun Song1, Bowen Li1, Ying Shen3thanks:   Corresponding Author., Zhi Wang1222FLOPS (Floating-Point Operations Per Second) measures the number of floating-point arithmetic operations (e.g., addition, subtraction, multiplication, division) that a computational unit can perform in one second. 1Shenzhen International Graduate School, Tsinghua University 2Shenzhen Technology University 3China Central Depository & Clearing Co., Ltd. donght24@mails.tsinghua.edu.cn, jiangjingyan@sztu.edu.cn wangzhi@sz.tsinghua.edu.cn
@@ 1 Introduction
The rapid advancement of LLMs has yielded remarkable progress across a wide range of domains DeepSeek-AI et al. (2025b). With model scales expanding from GPT-3’s Brown et al. (2020) 175 billion to DeepSeek-R1’s DeepSeek-AI et al. (2025a) 660 billion parameters, the computing resource demands of training LLMs have increased dramatically Jiang et al. (2024).
However, this exponential growth in computational requirements poses significant challenges. For individual researchers and small laboratories with limited resources, the demands are particularly prohibitive. Even for well-resourced organizations, confining LLM training within a single AI cluster faces challenges like: geographically distributed services required for latency optimization McMahan et al. (2017), inherent hardware bottlenecks limiting single-cluster scalability Athlur et al. (2022); Grattafiori et al. (2024), economic patterns requiring placement adaptation Liu et al. (2023), etc.
@@
Duan et al. (2024)
Efficient LLM Training, Centralized Infrastructures
Khan et al. (2023)
Decentralized Machine Learning, Geo-Distributed Machine Learning
Woisetschläger et al. (2024)
Federated Learning, Efficient Foundation Model Training
Ours
Decentralized LLM Training, Decentralized Infrastructures
@@
This survey systematically investigates challenges and solutions for decentralized LLM training. Compared to prior surveys on other distributed paradigms, our paper centers around LLM training with decentralized resources, as shown in Table 1. We aim to characterize the utilization of decentralized resources and analyze optimization methods in decentralized LLM training, thereby exploring potential research opportunities. Figure 1 illustrates the utilization paradigms of resources in decentralized LLM training and optimization objectives we categorize in this paper. To the best of our knowledge, our survey is the first to study recent advances in decentralized LLM training. Our work contains three distinct contributions:
@@
  • Up-to-date case studies and prospective future research directions. We compare two LLMs trained under different decentralized paradigms: one leveraging organizational resources and the other utilizing scattered, fragmented resources. We also propose potential directions for future research.
Figure 1: Decentralized LLM training paradigms of both community-driven and organizational decentralization. Community-driven decentralization utilizes pooling resources from communities with independent entities. These pooled resources can include local servers from researchers, instances from cloud services or volunteer communities, etc. Organizational decentralization involves consolidating resources from multiple clusters or even multiple datacenters managed by well-resourced organizations (e.g., technology giants and governments).
@@ 2.1 Geo-ML and FL
Geo-ML and FL both tackle challenges from data decentralization, yet each approach has its distinct focus. Geo-ML mainly tackles service latency and regulatory requirements by optimizing training processes across datacenters, leveraging hierarchical network topologies (e.g., high-bandwidth local area networks (LANs) within datacenters and limited-bandwidth WANs between datacenters) to utilize decentralized data and computational resources efficiently Hsieh et al. (2017). In contrast, FL focuses more on privacy protection McMahan et al. (2017), employing either centralized parameter-server or decentralized methods Lalitha et al. (2019); Xing et al. (2021), with resources from decentralized clusters or edge devices.
@@ 2.2 De-ML
De-ML has a dual meaning in terms of decentralization. From the distributed strategy perspective, it mitigates communication bottlenecks inherent in parameter server architectures through peer-to-peer networking Hegedűs et al. (2019); Warnat-Herresthal et al. (2021). From the resource utilization perspective, it offers a cost-effective and flexible strategy that maximizes the utilization of geographically dispersed computing resources Yuan et al. (2022). While both De-ML and Geo-ML involve training architectures, they differ fundamentally in their approach—the former adopts peer-to-peer topology where nodes communicate directly, while the latter relies on hierarchical centralized architecture both across and within datacenters.
@@ 2.4 Decentralized LLM Training
[…]
@@ 3.1 Communication Constraint
In distributed training, computational devices need to communicate frequently. However, decentralized resources have more limited communication capabilities compared to centralized ones.
Community-driven decentralized resources, combining with local servers and cloud instances, communicate through LANs or even WANs during model training with bandwidths under 10 Gb/s Azurespeed (2024). In contrast, organization-driven decentralized training can reach over 500 GB/s of bandwidth within a server by NVLink, or over 200 Gb/s within a cluster using InfiniBand. However, when the scale of resources expands across multiple geographically distributed clusters or datacenters, the bandwidth can drop to only hundreds of Mb/s Xiang et al. (2022). To alleviate communication bottlenecks, strategies such as gradient compression Lu et al. (2024b, 2025b) and delayed aggregation Zhu et al. (2021); Lu et al. (2025a) are often employed during distributed training.
@@ 3.3 Cost Volatility
However, reducing costs with decentralized resources is often more challenging than centralized setups. In community-driven decentralization, fragmented resources such as preemptive cloud instances are relatively cheaper, but their prices are highly volatile Lee and Son (2017). Additionally, resource instability can lead to re-computation and inter-node waiting, prolonging training process and increasing overall costs. For organizational decentralization, accurately modeling energy consumption and optimizing resource scheduling become increasingly complex, owing to the demand for massive resources and the inherently cross-datacenter nature of this paradigm Faiz et al. (2024); Choudhury et al. (2024).
@@ 4 Community-Driven Decentralization
Training LLMs under community-driven paradigm often encounters challenges, including hardware heterogeneity, limited communication bandwidth, resource instability, and economic volatility. This chapter focuses on these challenges and provides an in-depth discussion of related research efforts. We summarize and compare selected studies in Table 2. A more comprehensive paper list is presented in Figure 3 of Appendix B.
@@
DeDloc Diskin et al. (2021) DP with extremely large batches BERT (170M) 91 devices (RTX 2060, K80, etc.)
@@ 4.1 Communication Optimization
In PP strategy, gradients are exchanged within a stage, while both gradients and activations are transmitted across stages. Varuna Athlur et al. (2022) leverages rule-based policy to adjust PP depth based on the available GPU count to better accommodate bandwidth constraints. To allocating tasklets requiring high communication volume to computing units with faster connections, DTFM Yuan et al. (2022) uses a two-level approach: a balanced graph partitioning problem for DP within each stage, and a joint graph matching and traveling salesman problem for the entire PP process. More dynamically, SWARM Ryabinin et al. (2023) optimizes the PP process by enabling real-time adjustments during each iteration. In this strategy, each PP stage uses multiple candidate devices. When a device outperforms others, it processes inputs from multiple slower predecessors and distributes outputs to multiple slow successors, maximizing bandwidth utilization.
@@ 4.2 Heterogeneity Awareness
Bandwidth constraints necessitate optimization of communication volume in decentralized LLM training, while network heterogeneity introduces additional challenges in communication scheduling.
To deal with network heterogeneity, BACombo Jiang et al. (2020) uses a bandwidth-aware worker selection strategy, enabling efficient gradient splitting and scheduling across multiple connections. DTFM Yuan et al. (2022) employs a communication matrix to model bandwidth differences, and it solves the hierarchical optimization problem to minimize communication cost. In contrast to static scheduling, DeDLOC Diskin et al. (2021) dynamically adapts its strategy by switching between All-Reduce, parameter servers, and decentralized SGD based on real-time network conditions. To further optimize communication paths, NETSTOR Li et al. (2024) employs a multi-root adaptive synchronization topology to dynamically allocate tasks based on bandwidth and balancing communication loads with auxiliary paths. Additionally, hardware characteristics can be integrated into the scheduling process. Holmes Yang et al. (2024) automatically selects heterogeneous network interface cards (NICs, including InfiniBand, RoCE, Ethernet) for diverse distributed strategies (DP, PP and TP) across heterogeneous clusters.
@@ Discussion.
The effective utilization of decentralized resources is fundamentally constrained by the ability to manage system heterogeneity. Current frameworks rarely account for cross-architecture collaboration (e.g., NVIDIA GPUs with CUDA [^3] https://developer.nvidia.com/cuda-toolkit and Huawei Ascend GPUs with CANN [^4] https://www.hiascend.com/software/cann), which limits their potential. Furthermore, existing heterogeneity-aware strategies mainly focus on protocol or topology optimization, overlooking critical semantic aspects of LLM training dynamics, such as layer sensitivity and gradient rank Refael et al. (2025). Developing semantics-guided, architecture-agnostic frameworks may enhance scalability and resource efficiency in decentralized LLM training.
@@ 4.4 Economics
Additionally, several studies propose analytical cost models to guide deployment decisions. Strati et al. (2024) develops a training cost estimator that highlights the benefits of intra-region communication for improving throughput and minimizing expenses. Phalak et al. (2024) extends this to a multi-cloud, multi-geography scenario, incorporating serverless compute and VM selection into a unified model for performance-cost optimization.
@@ 5 Organizational Decentralization
In organizational paradigms, besides the optimizations involved in community-driven decentralization, well-resourced organizations consider objectives that extend beyond individual systems to include datacenter-level enhancements, such as energy efficiency, system throughput [^5] System throughput refers to the amount of data or tasks a training system can process per unit of time, reflecting its overall processing capacity and efficiency. within and across datacenters. Representative related papers are compactly summarized in Figure 2.
Organizational
Decentralization
Energy Efficiency §5.1 TowardDC Wang et al. (2023), CAFE Bian et al. (2024), LLMCarbon Faiz et al. (2024), CAFTM Park et al. (2024), CloudSimPer Song et al. (2024). System Throughput §5.2 Yugong Huang et al. (2019b), TFAS Fan et al. (2024), MAST Choudhury et al. (2024), ACME Hu et al. (2024).
@@ 5.1 Energy Efficiency
Training LLMs within datacenters entails substantial energy consumption, making carbon efficiency a critical concern for large organizations. For a single job, predicting the carbon footprint of current training methods for large models can help optimize training strategies, thereby reducing carbon emissions Faiz et al. (2024). Resource selection across datacenters also plays a vital role, enabling a trade-off between model performance and environmental impact Bian et al. (2024) At the datacenter level, implementing job scheduling strategies that are aware of carbon emissions can further decrease the overall carbon footprint Park et al. (2024); Song et al. (2024). In addition, digital twin technologies hold significant potential for enabling real-time monitoring, control, and optimization of datacenter operations, thereby enhancing energy efficiency during LLM training Wang et al. (2023).
@@ 5.2 System Throughput
When training large models across datacenters, it is crucial to design resource scheduling policies that account for the characteristics of training processes, such as 4D parallelism and synchronization requirements, as well as the infrastructure characteristics including low-bandwidth inter-datacenter communication Fan et al. (2024); Hu et al. (2024). Traditional cross-datacenter schedulers often fall short, as they struggle to adapt to the dynamic and resource-intensive nature of LLM training workloads Huang et al. (2019b). To address these challenges, MAST Choudhury et al. (2024), a global scheduling system, effectively orchestrates the training of LLama-3 Grattafiori et al. (2024) across 16,000 H100 GPUs, achieving both load balancing and fault tolerance across datacenters.
@@ 6 Case Study
This section examines two representative models, Llama-3 and INTELLECT-1, which we considered as examples of organizational and community-driven decentralization, respectively. Comparative configurations are listed in Table 3.
@@
Llama-3, an open-source foundation model family from Meta, scales up to 405 billion parameters Grattafiori et al. (2024). It is trained on 15 T multilingual tokens, requiring a significant amount of FLOPs for computation. To achieve this, Llama-3 utilizes 16,000 H100 GPUs with a global-scale scheduler Choudhury et al. (2024) and 4D parallel training (as illustrated in Appendix A), achieving 38-43% Model Fractional Utilization (MFU). Storage is managed by the Tectonic distributed file system Pan et al. (2021), offering 240 PB capacity and optimized throughput to reduce GPU idle time. For networking, Llama-3 leverages RoCE and InfiniBand within a three-layer topology Lee et al. (2024); Gangidi et al. (2024), enhanced by load balancing and congestion control for efficient communication across 24,000 GPUs. As a representative open-source LLM, Llama-3 leverages interconnected clusters orchestrated by a global-scale scheduler, providing critical insights for organizational decentralized LLM training paradigms.
@@
INTELLECT-1 Jaghouar et al. (2024a) , an open-source 10-billion-parameter LLM trained with decentralized resources. It employs hierarchical parameter aggregation and int8 quantization to minimize bandwidth usage, while VPN is integrated to ensure stability in low-bandwidth networks. For fault tolerance, INTELLECT-1 utilizes ElasticDeviceMesh for node management and phased checkpointing to minimize training interruptions, with peer-to-peer transfer enabling rapid checkpoint recovery. To optimize memory utilization, INTELLECT-1 integrates FSDP2 Zhao et al. (2023) and CPU offloading, enhancing the system efficiency and scalability. Furthermore, Intellect-2 Team et al. (2025), the successor model of INTELLECT-1, extends the capabilities of decentralized LLMs through decentralized post-training with reinforce learning. This community-driven decentralized paradigm democratizes AI model development, preventing monopolization and fostering open-source innovation with decentralized resources.
@@
Effective Training Time $\geq 90\%$ 83%
Processed Tokens 15 T 1 T